# Bulk photovoltaic effect in ferroelectric and antiferroelectric phases of antimony sulphoiodide investigated by means of ab-initio simulations

### v1

@@ 3.3 Band structure
The band structures of the FE and AFE phases, calculated within the GGA and including SOC, are displayed in Fig. S2(a) and (b), respectively. The results for the FE phase are consistent with previous studies [65, 32, 73]. The bands are quite wide along paths in the Brillouin zone parallel to the $k_{z}$ direction (e.g., Z-$\Gamma$ line or X-U), while they narrow along the in-plane directions (e.g., $\Gamma$-X line), reflecting the weak inter-chain hybridization. Compared to the GGA, HSE gives a rigid upward shift of the conduction bands without significantly modifying the main electronic features and the overall band dispersion. The HSE band structure is shown in the SI.
@@
The main differences between the FE and AFE phases are observed in some bands that are degenerate in one phase but split by the SOC in the other. This is a direct consequence of the change in the crystal space group during the phase transition. A particularly insightful analysis can be made around the $\Gamma$ point; despite the intricate band structure, one can resort to an effective linear spin-momentum Hamiltonian for a qualitative understanding. In the FE phase, the point group symmetry at $\Gamma$ is $C_{2v}$. Thus, the effective SOC terms acquire the form $\alpha_{xy}k_{x}\tau_{y}$ and $\alpha_{yx}k_{y}\tau_{x}$, where $\tau_{x}$ and $\tau_{y}$ are the Pauli matrices, and $\alpha_{xy}$ and $\alpha_{yx}$ are coupling parameters representing the non-zero components of a rank-two spin-momentum coupling pseudotensor [74]. These effective SOC terms can be-rewritten in the standard Rashba and Dresselhaus forms by performing an appropriate unitary transformation and lead to a characteristic in-plane spin-texture [65]. Because of this peculiar feature, SbSI is referred to as a FE Rashba-Dresselhaus semiconductor.
@@
In contrast, in the AFE phase, the point group symmetry at $\Gamma$ is $D_{2}$, leading to a different effective Hamiltonian. Specifically, the SOC terms now take the form $\alpha_{xx}k_{x}\tau_{x}$, $\alpha_{yy}k_{y}\tau_{y}$, and $\alpha_{zz}k_{z}\tau_{z}$, where $\alpha_{xx}$, $\alpha_{yy}$, $\alpha_{zz}$ are the non-zero components of a spin-momentum coupling pseudotensor consistent with the $D_{2}$ symmetry. These SOC terms give rise to a Dresselhaus-Weyl spin-texture [75]. The effect of the $\alpha_{zz}k_{z}\tau_{z}$ term is particularly evident in the band structure, where bands along the Z-$\Gamma$ direction, which are two-fold degenerate in the FE phase, become spin-split in the AFE phase.
@@
The spin-textures obtained from the DFT calculations are shown in Fig. 4, with the blue arrows representing the directions of expectation values of Pauli spin-matrices. In particular, panel (a) shows the plot in the $k_{z}=0$ plane at the constant energy surface $E-E_{\mathrm{VBM}}$=1.75 eV above the valence band maximum (VBM) (see the horizontal dashed line in Fig. S2) for the FE phase, while panel (b) shows the plot in the $k_{x}=0$ plane at the same constant energy surface for the AFE phase. Panels (c) and (d) are zoomed-in views of panels (a) and (b) around the $\Gamma$ point. In panel (c), we clearly observe the characteristic Rashba-Dresselhaus spin-texture, with the Dresselhaus component dominating, consistent with previous reports for the FE phase [65]. In contrast, panel (d) reveals that the spin-texture of the AFE phase can be approximated as $k_{y}\tau_{y}-k_{z}\tau_{z}$, displaying a Weyl character as predicted by the symmetry analysis.
@@
The observed differences in the band structure and, even more, in the spin texture between the two phases reveal a particularly interesting effect. Since the FE and AFE phases are energetically competitive, it might be possible to explore means to stabilize one phase over the other, thereby tuning the effective spin-charge coupling and, consequently, the spintronic properties of the material. In the AFE phase, in particular, the Weyl-like spin texture is expected to give rise to the so-called ”unconventional” current-induced spin polarization [76, 77, 78], where a charge current induces spin polarization of the conduction electrons along the current direction. This phenomenon, recently reported in Te, has drawn significant attention for the design of spintronic devices [76]. In SbSI, it can be further combined with the BPVE, investigated in the following, resulting in the integration of multiple functionalities.
@@ 3.4 Bulk photovoltaic effect
$j^{a}=2\sum_{bc}{\sigma^{abc}(\omega})\mathrm{Re}[E^{b}(\omega)E^{c}(-\omega)]+t\sum_{bc}{\eta^{abc}(\omega})\mathrm{Im}[E^{b}(\omega)E^{c}(-\omega)]$ (1)
where the first term represents the shift current (linear photogalvanic effect) and the second term corresponds to the injection current (circular photogalvanic effect). Both currents are expected to appear in both the FE and AFE phases, due to their non-centrosymmetric space groups, and more specifically, to their gyrotropy. $\sigma^{abc}(\omega)$ denotes a component of the shift photoconductivity tensor, which is real, while $\eta^{abc}(\omega)$ denotes a component of the circular injection photoconductivity tensor, which is imaginary. They are defined in the SI following Refs. [11, 12]. The indices $a,b,$ and $c$ label vector and tensor components; in particular, $a$ indicates the direction of the photocurrent. $t$ represents the electron relaxation time. The shift current corresponds to the dc part of the second-order response to the dipole electric field of light. In contrast, the injection photoconductivity corresponds to the rate of change of the interband dc part of the second-order response.
The photoconductivity tensors presented in the following are all computed starting from a Wannier interpolation of the GGA band structure calculated with SOC, as detailed in the Methods section and in the SI. The results obtained from the HSE band structure are very similar to the GGA ones, except for the position of the photoconductivity onset energy, which is shifted due to the (roughly rigid) upward shift of the conduction bands and related increase of the band gap. Due to this uncertainty on the band gap size, in the following, we will present all the results with the onset energy aligned with the experimental band gap value, as for common practice in the literature [11]. Since the value of the band gap in the AFE phase is similar to that in the FE phase, we shift the photocurrents in AFE configuration by the same amount.
@@ 3.4.1 Linear photovoltaic effect
In the FE phase, the shift photoconductivity tensor of SbSI has five independent non-zero components, namely $\sigma^{zzz}$, $\sigma^{zxx}$, $\sigma^{zyy}$, $\sigma^{xxz}$ and $\sigma^{yyz}$, as allowed by the system’s space group, Pna21. They are displayed in Fig. 5(a) as a function of the energy. All other components can be obtained from these independent components by exchanging the last two indices, related to the light polarization. The largest component is $\sigma^{zzz}$, corresponding to the electric field being polarized along the $c$ axis and the current flowing in the same direction. The shift current is thus larger when the incident light is parallel to the FE distortion direction compared to the case when it is perpendicular. In the visible energy region ($E\sim 1.6-3.3$ eV), $\sigma^{zzz}$ displays two prominent peaks, reaching about 30 $\mu$A/V2 at 2 eV and 47 $\mu$A/V2 at 3 eV. These peak values are significantly higher than those for prototypical FE compounds, like for BiFeO3 and BaTiO3, where the maximum photoconductivities in the visible energy range are 0.05 $\mu$A/V2 and 5 $\mu$A/V2, respectively [30]. Hence, SbSI is potentially an excellent material candidate for bulk photovoltaic conversion.
@@
When transitioning to the AFE phase, the shift photoconductivity tensor is found to change compared to the FE case acquiring different non-zero components, as a result of the change in the crystal space group (see SI). Specifically, the independent components become $\sigma^{xyz}$, $\sigma^{yxz}$ and $\sigma^{zxy}$. They are displayed in Fig. 5(b). The largest one is $\sigma^{xyz}$. Thus, the shift current in the AFE phase is expected to be maximized in the $x$ direction, which is perpendicular to the chains, conversely to what is predicted in the FE phase. $\sigma^{xyz}$ presents peaks of approximately 17 $\mu$A/V2 near 2 eV and 20 $\mu$A/V2 at 3 eV. These photoconductivity values are lower than those in the FE phase at similar energies, but still large if compared to the response of the prototypical bulk FE oxides [30].
@@ 3.4.2 One-chain system
The one-chain system is modelled considering the FE unit cell, removing one of the two chains, and increasing the lattice parameters along the transverse directions to $a=b=12$, to avoid interaction between periodically repeated images. Then, following Refs. [11, 36], the shift photoconductivity tensor components for the single chain $\sigma^{abc}_{SC}$ are estimated by rescaling the results from the 3D supercell as $\sigma^{abc}_{SC}=\frac{a}{w_{x}}\frac{b}{w_{y}}\sigma^{abc}$ where $w_{a}$ and $w_{b}$ are the widths of the chain along the $x$ and $y$ directions, taken approximately equal to 2.62 Å and 6.57 Å (see SI).
The results for the one-chain system are displayed in Fig. 6, where only the components $\sigma^{zzz}$ and $\sigma^{zyy}$ are non-zero. In particular, $\sigma^{zzz}$ reaches peak values one order of magnitude larger than those for the bulk systems, especially in the energy range between 2.5 and 3.2 eV. The shape of the spectrum is characterized by sharp peaks which can be clearly related to the van Hove singularities in the DOS, shown in the SI. This confirms the significant role these singularities play in enhancing the shift photoconductivity, similar to what has been reported in monolayer ferroelectrics.
@@ 3.4.3 Circular photogalvanic effect
In the FE state, the circular injection photoconductivity tensor has two independent components, namely $\eta^{xxz}$ and $\eta^{yyz}$, allowed by symmetry. All other non-zero components are obtained from these by exchanging the last two indices and reversing the sign (see SI). Among these two independent components, $\eta^{yyz}$ is considerably larger than $\eta^{xxz}$ in the visible light energy range. It presents a first small peak of approximately 4 $\times$ 108 A/V2s at around 2.1 eV, followed by a significantly larger peak of 14 $\times$ 108 A/V2s at around 2.4 eV, and a sharp negative peak reaching -19 $\times$ 108 A/V2s at about 2.8 eV. As the energy increases, further large peaks are observed with the absolute value of the photoconductivity remaining quite large even outside the visible range.
@@
It is particularly important to note that all $\eta^{zbc}$ components in bulk FE SbSI are zero (for any $a$ and $b$ indices), indicating that the circular injection current is absent along the chains, but can only emerge in the transverse $x$ and $y$ directions, regardless of the light polarization. Interestingly, although the electronic bands are relatively flat in these perpendicular directions (i.e. the electron group velocity is quite small) owing to the weak inter-chain hybridization, the obtained photoconductivity peak values are an order of magnitude higher than those reported in other bulk materials, such as CdSe, CdS [26, 79, 80] and TaIrTe4 [12], which have been recently studied by DFT calculations. This indicates that simple arguments based on qualitative inspections of the band structure do not always hold in the analysis of photocurrents. Accurate quantum mechanical calculations are generally required to obtain reliable predictions.
@@
Turning to the AFE phase, we find that the independent components of the circular injection photoconductivity become $\eta^{xyz}$, $\eta^{yxz}$ and $\eta^{zxy}$. Thus, at variance with what happens in the FE phase, the circular injection current can be measured along all directions, including along the chains (longitudinal direction), for an opportune choice of the light polarization.
@@
The important result that the circular photogalvanic effect generates only transverse photocurrents in the FE phase, but also longitudinal photocurrents in the AFM phase, can be further confirmed through an alternative analysis. Specifically, by employing a distinct yet fully equivalent definition to Eq. (1), the circular photogalvanic effect is often characterized using a rank-two pseudotensor with components $\gamma_{ab}$ [74], derived from the circular injection photoconductivity as $\gamma_{ab}\propto-i\epsilon_{bcd}\eta^{acd}$ (note the convention that repeated indices are summed over) [81]. This so-called circular photogalvanic pseudotensor has the same structure as the optical activity tensor, and its components transform like any other quantities that can be expressed as components of a rank-two pseudotensor [82]. These include the effective SOC terms introduced in section 3.3, such that a non-zero $\alpha_{ab}$ parameter must correspond a non-zero $\gamma_{ab}$. In other words, we can also predict the character of the circular photogalvanic effect by analyzing the spin texture, and vice versa [74]. In the case of FE SbSI, we find that the only two non-zero components of the circular photogalvanic pseudotensor are $\gamma_{xy}$ and $\gamma_{yx}$, consistent with the non-zero spin-momentum coupling parameters $\alpha_{xy}$ and $\alpha_{yx}$, describing the Rashba-Dresselhaus in-plane spin-texture in the band structure. In contrast, in the AFE phase, there are three non-zero and independent components, $\gamma_{xx}$, $\gamma_{yy}$, and moreover, $\gamma_{zz}$, which accounts for the longitudinal photocurrent. These components correspond to the effective spin-momentum coupling parameters $\alpha_{xx}$, $\alpha_{yy}$, and $\alpha_{zz}$ identified in the band structure analysis and are consistent with the Weyl-Dresselhaus spin-texture.
@@
In summary, our results show that the circular photovoltaic effect can be used to probe the FE-AFE transition in the same way as the linear photovoltaic effect since the non-zero components of the photoconductivity change with the phase symmetry. However, measuring the circular injection currents can provide additional information about spin-dependent properties of the material. This feature may eventually be important for the integration of the BPVE with spintronic functionalities.

### v2

@@ 3.3 Band structure
The band structures of the FE and AFE phases, calculated within the GGA and including SOC, are displayed in Fig. S3(a) and (b), respectively. The results for the FE phase are consistent with previous studies [79, 45, 90]. The bands are quite wide along paths in the Brillouin zone parallel to the $k_{z}$ direction (e.g., Z-$\Gamma$ line or X-U), while they narrow along the in-plane directions (e.g., $\Gamma$-X line), reflecting the weak inter-chain hybridization. Compared to the GGA, HSE gives a rigid upward shift of the conduction bands without significantly modifying the main electronic features and the overall band dispersion. The HSE band structure is shown in Section S5 of the SM.
@@
The main differences between the FE and AFE phases are observed in some bands that are degenerate in one phase but split by the SOC in the other. This is a direct consequence of the change in the crystal space group during the phase transition. A particularly insightful analysis can be made around the $\Gamma$ point; despite the intricate band structure, one can resort to an effective linear spin-momentum Hamiltonian for a qualitative understanding. In the FE phase, the point group symmetry at $\Gamma$ is $C_{2v}$. Thus, the effective SOC terms acquire the form $\alpha_{xy}k_{x}\tau_{y}$ and $\alpha_{yx}k_{y}\tau_{x}$, where $\tau_{x}$ and $\tau_{y}$ are the Pauli spin matrices, and $\alpha_{xy}$ and $\alpha_{yx}$ are coupling parameters representing the non-zero components of a rank-two spin-momentum coupling pseudotensor [91]. These effective SOC terms can be-rewritten in the standard Rashba and Dresselhaus forms by performing an appropriate unitary transformation and lead to a characteristic in-plane spin texture [79]. Because of this peculiar feature, SbSI is referred to as a FE Rashba-Dresselhaus semiconductor.
@@
In contrast, in the AFE phase, the point group symmetry at $\Gamma$ is $D_{2}$, leading to a different effective Hamiltonian, corresponding to a spin texture with spin components that can also point outside the plane, i.e. along the $z$ axis. The effective SOC terms now take the form $\alpha_{xx}k_{x}\tau_{x}$, $\alpha_{yy}k_{y}\tau_{y}$, and $\alpha_{zz}k_{z}\tau_{z}$, where $\alpha_{xx}$, $\alpha_{yy}$, $\alpha_{zz}$ are the non-zero components of a spin-momentum coupling pseudotensor consistent with the $D_{2}$ symmetry. These SOC terms give rise to a Dresselhaus-Weyl spin texture [92]. The effect of the $\alpha_{zz}k_{z}\tau_{z}$ term is particularly evident in the band structure, where bands along the Z-$\Gamma$ direction, which are two-fold degenerate in the FE phase, become spin-split in the AFE phase.
@@
The spin textures obtained from the DFT calculations are shown in Fig. 4, with the blue arrows representing the directions of the expectation values of the Pauli spin matrices. In particular, panel (a) shows the plot in the $k_{z}=0$ plane at the constant energy surface $E-E_{\mathrm{VBM}}=1.75$ eV above the valence band maximum (VBM) (see the horizontal dashed line in Fig. S3) for the FE phase, while panel (b) shows the plot in the $k_{x}=0$ plane at the same constant energy surface for the AFE phase. Panels (c) and (d) are zoomed-in views of panels (a) and (b) around the $\Gamma$ point. In panel (c), we clearly observe the characteristic Rashba-Dresselhaus spin texture, consistent with previous reports for the FE phase [79]. In contrast, panel (d) shows, in the AFM phase, the emergence of a spin-z polarization, which is absent in the FE phase.
@@
The observed differences in the band structure and, even more, in the spin texture between the two phases reveal a particularly interesting effect. Since the FE and AFE phases are energetically competitive, it might be possible to explore means to stabilize one phase over the other, thereby tuning the effective spin-charge coupling and, consequently, the spintronic properties of the material. In the AFE phase, in particular, the spin texture with $\sim k_{z}\tau_{z}$ character is expected to give rise to the so-called “unconventional” current-induced spin polarization [93, 94, 95], where a charge current induces spin polarization of the conduction electrons along the current direction. This phenomenon, recently reported in Te, has drawn significant attention for the design of spintronic devices [93]. In SbSI, it can be further combined with the BPVE, investigated in the following, resulting in the integration of multiple functionalities.
@@ 3.4 Bulk photovoltaic effect
$j^{a}$ $=2\sum_{bc}\sigma^{abc}(\omega)\,\mathrm{Re}[E^{b}(\omega)E^{c}(-\omega)]$
$\quad+t\sum_{bc}\eta^{abc}(\omega)\,\mathrm{Im}[E^{b}(\omega)E^{c}(-\omega)]$ (1)
where the first term represents the shift current (linear photogalvanic effect) and the second term corresponds to the injection current (circular photogalvanic effect). Both currents are expected to appear in both the FE and AFE phases, due to their non-centrosymmetric space groups, and more specifically, to their gyrotropy. $\sigma^{abc}(\omega)$ denotes a component of the shift photoconductivity tensor, which is real, while $\eta^{abc}(\omega)$ denotes a component of the circular injection photoconductivity tensor, which is imaginary. They are defined in Section S7 of the SM following Refs. [[13]] and [[14]]. The indices $a,b,$ and $c$ label vector and tensor components; in particular, $a$ indicates the direction of the photocurrent. $t$ represents the electron relaxation time. The shift current corresponds to the dc part of the second-order response to the dipole electric field of light. In contrast, the injection photoconductivity corresponds to the rate of change of the interband dc part of the second-order response.
The photoconductivity tensors presented in the following are all computed starting from a Wannier interpolation of the GGA band structure calculated with SOC, as detailed in Section 2. The results obtained from the HSE band structure are very similar to the GGA ones, except for the position of the photoconductivity onset energy, which is shifted due to the (roughly rigid) upward shift of the conduction bands and related increase of the band gap. Due to this uncertainty on the band gap size, in the following, we will present all the results with the onset energy aligned with the experimental band gap value, as for common practice in the literature [13]. Since the value of the band gap in the AFE phase is similar to that in the FE phase, we shift the photocurrents in AFE configuration by the same amount.
@@ 3.4.1 Linear photovoltaic effect
In the FE phase, the shift photoconductivity tensor of SbSI has five independent non-zero components, namely $\sigma^{zzz}$, $\sigma^{zxx}$, $\sigma^{zyy}$, $\sigma^{xxz}$ and $\sigma^{yyz}$, as allowed by the system’s space group, Pna21. They are displayed in Fig. 5(a) as a function of the energy, $E=\hbar\omega$. All other components can be obtained from these independent components by exchanging the last two indices, related to the light polarization. The largest component is $\sigma^{zzz}$, corresponding to the electric field being polarized along the $c$ axis and the current flowing in the same direction. The shift current is thus larger when the incident light is parallel to the FE distortion direction compared to the case when it is perpendicular. In the visible energy region ($E\sim 1.6-3.3$ eV), $\sigma^{zzz}$ displays two prominent peaks, reaching about 30 $\mu$A/V2 at 2 eV and 47 $\mu$A/V2 at 3 eV. These peak values are significantly higher than those for prototypical FE compounds, like for BiFeO3 and BaTiO3, where the maximum photoconductivities in the visible energy range are 0.05 $\mu$A/V2 and 5 $\mu$A/V2, respectively [43]. Hence, SbSI is potentially an excellent material candidate for bulk photovoltaic conversion.
@@
When transitioning to the AFE phase, the shift photoconductivity tensor is found to change compared to the FE case acquiring different non-zero components, as a result of the change in the crystal space group (see SM). Specifically, the independent components become $\sigma^{xyz}$, $\sigma^{yxz}$ and $\sigma^{zxy}$. They are displayed in Fig. 5(b). The largest one is $\sigma^{xyz}$. Thus, the shift current in the AFE phase is expected to be maximized in the $x$ direction, which is perpendicular to the chains, conversely to what is predicted in the FE phase. $\sigma^{xyz}$ presents peaks of approximately 17 $\mu$A/V2 near 2 eV and 20 $\mu$A/V2 at 3 eV. These photoconductivity values are lower than those in the FE phase at similar energies, but still large if compared to the response of the prototypical bulk FE oxides [43].
@@ 3.4.2 One-chain system
The one-chain system is modelled considering the FE unit cell, removing one of the two chains, and increasing the lattice parameters along the transverse directions to $a=b=12$, to avoid interaction between periodically repeated images. Then, following Refs. [[13]] and [[48]], the shift photoconductivity tensor components for the single chain $\sigma^{abc}_{SC}$ are estimated by rescaling the results from the 3D supercell as $\sigma^{abc}_{SC}=\frac{a}{w_{x}}\frac{b}{w_{y}}\sigma^{abc}$ where $w_{a}$ and $w_{b}$ are the widths of the chain along the $x$ and $y$ directions, taken approximately equal to 2.62 Å and 6.57 Å (see SM).
The results for the one-chain system are displayed in Fig. 6, where only the components $\sigma^{zzz}$ and $\sigma^{zyy}$ are non-zero. In particular, $\sigma^{zzz}$ reaches peak values one order of magnitude larger than those for the bulk systems, especially in the energy range between 2.5 and 3.2 eV. The shape of the spectrum is characterized by sharp peaks which can be clearly related to the van Hove singularities in the DOS, which is shown in Section S6 of the SM. This results confirms the significant role these singularities play in enhancing the shift photoconductivity, similar to what has been reported in monolayer ferroelectrics.
@@ 3.4.3 Circular photogalvanic effect
In the FE state, the circular injection photoconductivity tensor has two independent components, namely $\eta^{xxz}$ and $\eta^{yyz}$, allowed by symmetry, as shown in Fig. 7(a). All other non-zero components are obtained from these by exchanging the last two indices and reversing the sign (see Section S7-C-2 in the SM). Among these two independent components, $\eta^{yyz}$ is considerably larger than $\eta^{xxz}$ in the visible light energy range. It presents a first small peak of approximately 4 $\times$ 108 A/V2s at around 2.1 eV, followed by a significantly larger peak of 14 $\times$ 108 A/V2s at around 2.4 eV, and a sharp negative peak reaching -19 $\times$ 108 A/V2s at about 2.8 eV. As the energy increases, further large peaks are observed with the absolute value of the photoconductivity remaining quite large even outside the visible range.
@@
It is particularly important to note that all $\eta^{zbc}$ components in bulk FE SbSI are zero (for any $a$ and $b$ indices), indicating that the circular injection current is absent along the chains, but can only emerge in the transverse $x$ and $y$ directions, regardless of the light polarization. Interestingly, although the electronic bands are relatively flat in these perpendicular directions (i.e., the electron group velocity is quite small) owing to the weak inter-chain hybridization, the obtained photoconductivity peak values are an order of magnitude higher than those reported in other bulk materials, such as CdSe, CdS [28, 96, 97] and TaIrTe4 [14], which have been recently studied by DFT calculations. This indicates that simple arguments based on qualitative inspections of the band structure do not always hold in the analysis of photocurrents. Accurate quantum mechanical calculations are generally required to obtain reliable predictions.
@@
Turning to the AFE phase, we find that the independent components of the circular injection photoconductivity become $\eta^{xyz}$, $\eta^{yxz}$ and $\eta^{zxy}$, as shown in Fig. 7(b). Thus, at variance with what happens in the FE phase, the circular injection current can be measured along all directions, including along the chains (longitudinal direction), for an opportune choice of the light polarization.
@@
The important result that the circular photogalvanic effect generates only transverse photocurrents in the FE phase, while both transverse and longitudinal photocurrents appear in the AFE phase, can be further confirmed through an alternative analysis. Specifically, the rank-three circular injection photoconductivity tensor can be recast into a rank-two pseudotensor of components $\gamma_{ab}\propto-i\epsilon_{bcd}\eta^{acd}$ (see Ref. [98]). This so-called circular photogalvanic pseudotensor has the same structure as the optical activity tensor, and it transforms like any other rank-two pseudotensor [91, 99], including the effective SOC pseudotensor introduced in Section 3.3, and whose components are the spin-momentum coupling parameters. Therefore, the form of the photogalvanic pseudotensor must mirror that of the spin-momentum coupling and, in turn, reflects the underlying spin texture. In practice, this means that we can predict the nature of the circular photogalvanic effect by analyzing the spin texture — and vice versa [91].
In the case of FE SbSI, we find that the only two non-zero components of the circular photogalvanic pseudotensor are $\gamma_{xy}$ and $\gamma_{yx}$, which are consistent with the non-zero spin–momentum coupling parameters $\alpha_{xy}$ and $\alpha_{yx}$, describing the Rashba–Dresselhaus in-plane spin texture. In contrast, in the AFE phase, there are three non-zero and independent components: $\gamma_{xx}$, $\gamma_{yy}$, and $\gamma_{zz}$, the latter accounting for the longitudinal photocurrent. These components correspond to the effective spin–momentum coupling parameters $\alpha_{xx}$, $\alpha_{yy}$, and $\alpha_{zz}$ identified in the band structure analysis, and are consistent with the emergence of spin-$z$ polarization alongside the in-plane spin texture.
@@
In summary, our results show that the circular photovoltaic effect can be used to probe the FE-AFE transition in a similar way to the linear photovoltaic effect, as the non-zero components of the photoconductivity change with the symmetry of the crystal structure. However, measuring circular injection currents can provide additional insights into the spin-dependent properties of the material. This feature could eventually be important for integrating the BPVE with spintronic functionalities.